1: A project has proposed investment is (Year 0 ) 320K, (Year 1) 150K, 100K, 60K, 0, 0, and (Year 6) 25K over these years respectively. The income is 0,0,200K, 225K, 250K, 175K, and 125K over these same corresponding years. Use this same data for Questions 25-28. The required rate of return is 14%. What is the Net inflow in Year 3? 175K 60K 225K 165K

2:

The required rate of return is 14%. What is the present value of Year 6 net inflow?

90.9K
34.6K
100K
45.5K

3:

The required rate of return is 14%. What is the net present value of the project?

-63.3K
21.2K
32.0K
15.3K

4: At what rate of return is the NPV approximately zero?

17.2
14
15.5
10.7
Answer #1

Given,

Year Investments, Cash outflow Profit, Cash inflow
n A B
0 320000 0
1 150000 0
2 100000 200000
3 60000 225000
4 0 250000
5 0 175000
6 25000 125000

Note: Investments are cash outflows

Profits are Cash inflows.

(a) 225K

Net flow for third year = Cash inflow(Year 3) - Cash Outflow(Year 3 = 225,000 - 0 = 225,000

(b) 45.5K

We know that

NPV = F / [ (1 + i)^n ]

NPV = Net present value

F = Future Cash flow

i = discount rate

n = number of years

For the Year 6, Future cash flow, F = Net Cash flow = Cash Inflow(Year 6) - Cash Outflow(Year 6) = 125,000 - 25,000 = 100,000

i = 14%

n = 6 years

Hence by using the formula mentioned above the required NPV of netflow for year 6 = 100,000 /[1+14/100]^6 = 45,558.65

(c) 21.2K

For calculating the NPV of the total project we need to calculate the net present value of the cash flow of each year as we did for year 6. The calculation and the tabulation are as shown below:

Year Investments,Cash outflow Profit, Cash inflow Net Cash flow Net present value
n A B F = B-A NPV = F / [ (1 + i)^n ]
0 320000 0 -320000 -320000.00
1 150000 0 -150000 -131578.95
2 100000 200000 100000 76946.75
3 60000 225000 165000 111370.30
4 0 250000 250000 148020.07
5 0 175000 175000 90889.52
6 25000 125000 100000 45558.65
Net present value of project = 21206.35

The net present value of the project is equal to the sum of the net present value of the project for all the years, which is calculated in the table above and equal to 293,358.7Please follow the notations for clear understanding.

(d) 15.5%

By calculating the NPV for all the above-mentioned interest rate as below in the table:

Year Investments,Cash outflow Profit, Cash inflow Net Cash flow Net present value @ 14% Net present value @ 17.2% Net present value @ 15.5% Net present value @ 10.7%
n A B F = B-A NPV = F / [ (1 + i)^n ] NPV = F / [ (1 + i)^n ] NPV = F / [ (1 + i)^n ] NPV = F / [ (1 + i)^n ]
0 320000 0 -320000 -320000.00 -320000.00 -320000.00 -320000.00
1 150000 0 -150000 -131578.95 -127986.35 -129870.13 -135501.36
2 100000 200000 100000 76946.75 72802.25 74961.11 81602.74
3 60000 225000 165000 111370.30 102494.63 107087.31 121630.10
4 0 250000 250000 148020.07 132504.17 140479.22 166475.19
5 0 175000 175000 90889.52 79140.72 85138.92 105268.87
6 25000 125000 100000 45558.65 38586.41 42121.91 54339.33
Net present value of project = 21206.35 -22458.17 -81.66 73814.88

Clearly for the interest rate 15,5% NPV of the project is closer to zero.